OpenStudy (anonymous):

Find two real numbers whose sum is 18 and whose product is a minimum.

hartnn (hartnn):

this belongs to calculus ?

OpenStudy (anonymous):

no. advanced algebra.

OpenStudy (anonymous):

x + y = 18 y = 18 - x

OpenStudy (anonymous):

etc.etc.etc...

hartnn (hartnn):

so can you now find the function to minimize only in terms of 'x' or 'y' ?

OpenStudy (anonymous):

no. i really dont have any idea.

OpenStudy (anonymous):

Hint: Substitute!

hartnn (hartnn):

f = xy x+y= 18 so, y =18-x put this in xy what do u get ?

OpenStudy (anonymous):

x(18-x)

hartnn (hartnn):

so, f (x) = 18x - x^2 do you know how to minimize this ?

OpenStudy (anonymous):

Convert to vertex form, that should make it easy to solve :)

OpenStudy (anonymous):

use am<=gm

hartnn (hartnn):

or use the fact that for \(ax^2+bx+c\) the miminum value of x is -b/a

OpenStudy (anonymous):

Many ways ^.^